# Towards Flexible, Scalable, and Adaptive Multi-Modal Conditioned Face Synthesis

### 26 Dec 2023 v1

@@ 3.2 Entropy-Aware Modal-Adaptive Modulation
To synthesize facial images of high fidelity given multi-modal conditions, it is essential to consider conditional entropies across modalities. For uni-modal synthesis, one can freely specify the control strength to obtain results of high quality and fidelity. However, in the multi-modal synthesis scenario, a fixed control strength across all modalities may not be appropriate. For some low control strength, the conditions of high condition entropy fail to provide sufficient influence on image synthesis process. Simply increasing control strength may cause the generated images to be over-fit to some condition of low condition entropy. As shown in Figure  5, an appropriate control strength for textual modality is too strong for mask modality. To address this issue, we devise a entropy-aware modal-adaptive modulation where before the output noise is altered by a fixed, user-defined control strength $w$, it is finely adjusted according to given conditions and current denoising state. We use a weighting network that processes the fused features $C_{m}$, the combination of given conditions and modal surrogates, and the U-Net feature $f_{u}$. This network performs average pooling on these features, concatenates the results, and passes them through several linear layers to produce a weighting vector $w$. The modal surrogates encapsulate modality-specific priors. When combined with the given conditions, they offer crucial guidance on noise adjustment. The inclusion of U-Net features supplies real-time insights $f_{u}$ into the denoising trajectory. In addition to predicting a single base noise map based on U-Net feature $f_{u}$ we predict multiple ($K$) noise maps, each potentially capturing different aspects of the features that could be added to the image at the current step of the diffusion process. We finally obtain the output noise as:
$\begin{split}n=\frac{1}{K}\sum_{k}(w_{k}(n_{k}-n_{b})+n_{b}).\end{split}$ (7)
We start with a base noise pattern $n_{b}$ and introduce adjustment from other noise maps $n_{k}$ of which the influence is scaled by $w_{k}$ according to the specific requirements of the multimodal conditioned image synthesis process. The noise is finely modulated according to modal-specific information, given conditions and current state, enabling well-informed and strategic decisions in each step of de-noising and ultimately leading to synthesis result of high quality and high fidelity to multi-modal conditions.
TediGAN [34]
Compose [20, 22]
CoDiff [13]
T2I-Adapter [21]
ControlNet [37]
This woman looks extremely young. There is no beard on her face
He is in his middle age. The face is covered with beard of medium length.
TediGAN [34]
Compose [20, 22]
CoDiff [13]
T2I-Adapter [21]
ControlNet [37]
@@
Method Training Data Type FID $\downarrow$ Text(%) $\uparrow$ Mask (%) $\uparrow$
TediGAN [34] Uni-modal 116.04 24.38 86.37
Compose [20, 22] Uni-modal 127.39 24.22 77.34
CoDiff [13] Multi-modal 122.51 24.37 85.94
ControlNet [37] Multi-modal 136.41 25.70 85.44
T2I-Adapter [21] Multi-modal 139.82 26.11 79.84
Ours Uni-modal 103.14 24.70 90.16
4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments on the Celeb-HQ dataset [15] with 30k high-quality face images. We utilized mask from [18] and textual descriptions from [14]. We compared our method, focusing on parsing mask and text modalities, for fair comparison with existing multi-modal methods following [13]. Our approach allowed to efficiently include additional modalities such as sketches [34], 3D face models [7], and low-resolution images [32], enriching our range of synthesis tasks.
Figure 7: Comparison of synthesis results across training methodologies. Pink arrows indicate image generation with mask condition, while blue arrows represent text involvement. Our uni-modal training with modal surrogate enables a flexible and scalable face synthesis framework. Entropy-aware Modality-Adaptive Modulation further enhances the fidelity to given conditions.
Implementation details. We implement our method based on Latent Diffusion [27], where the images $(3\times H\times W)$are first mapped to latent features $(3\times\frac{H}{8}\times\frac{W}{8})$. The modal encoder of the low-resolution image is the same as the first stage encoder of LDM [27]. The encoder of text and low-resolution image is fixed during training, while the encoders of other modalities are updated in training. The modal surrogate for text is initialized as the textual encoder feature of input human face and others are randomly initialized.
@@ 4.2 Comparison Analysis
Quantitative comparisons. In Table  1, we benchmark our method’s FID, Text, and Mask accuracy in multi-modal facial synthesis against leading techniques. TediGAN[34] yield impressive results by optimizing each condition and using a high-resolution generator StyleGAN. ControlNet [37] and T2I-Adpter [21] score well in text accuracy, leveraging advanced text encoders[25] and inheriting capabilities from Stable Diffusion[27]. Compose[22, 20] lags behind due to its simplistic fusion of uni-modal model noise, overlooking crucial modality-specific and inter-modal dynamics. Our method, employing uni-modal training with modal surrogates and entropy-aware adaptive noise modulation, achieves superior results by effectively utilizing and integrating multi-modal controls.
Qualitative comparisons. Figure 6 showcases that our methods produce high-quality facial images, faithfully adhering to specified conditions. While other methods struggle with aligning to details like beards or mask shapes, often resulting in poor quality, our approach excels. Our generated images accurately match text descriptions and mask layouts, demonstrating our method’s proficiency in leveraging each modality’s conditions to create precise, high-fidelity facial images. Please refer to our supplementary materials for more comparison results.
Method Task FID $\downarrow$ Text(%$\uparrow$ Mask (%$\uparrow$
Multi-modal training w/o modal surrogate Mask&Text to Image 124.98 24.28 85.44
Mask to Image - - -
Text to Image - - -
Uncondition to Image - - -
Multi-modal training with modal surrogate Mask&Text to Image 112.25 24.42 89.91
Mask to Image 108.69 - 90.06
Text to Image 195.94 23.18 -
Uncondition to Image 167.21 - -
Uni-modal training with modal surrogate Mask&Text to Image 113.29 24.28 89.72
Mask to Image 109.83 - 88.97
Text to Image 114.07 24.48 -
Uncondition to Image 127.26 - -
Uni-modal training with modal surrogate Noise Reweighting Mask&Text to Image 103.14 24.70 90.16
Mask to Image 112.91 - 90.31
Text to Image 110.32 24.78 -
Uncondition to Image 129.52 - -
Table 2: Ablation results of uni-modal training with modal surrogates and entropy-aware modal-adaptive modulation.
@@ 4.3 Ablation Analysis
We first validate our uni-modal training with modal surrogates for facilitating flexible and scalable face synthesis. Next, we examine the effectiveness of our entropy-aware modal-adaptive modulation in fully leveraging the control of varing multi-modal conditions.
Uni-modal training with modal surrogate. Our baseline is to directly train a multi-modal synthesis network using multi-modal data. As shown in the top row of Table  2 and Figure 7 (a), this methods effectively generates images combining mask and text inputs. However, it’s limited to specific mask-text combinations and doesn’t support unconditional or single-modality generation. Enhancing our initial approach, we integrated modal surrogates, enabling the network to combine condition with its corresponding modal surrogate for image generation. This method, shown in the second top-down row of Table  2 and Figure  7 (b), supports not only multi-modal synthesis but also uni-modal or unconditional generation, thanks to the modal surrogate’s representative capabilities. It is noticed a marked performance gap between text-to-image and mask-to-image synthesis: the FID text-to-image was 195.94, much higher than FID for mask-to-image 108.69. This disparity is largely due to the lower conditional entropy of masks, which offer clearer visual cues, lead the network overly rely on mask and thus neglect text. Figure  7 (b) shows that visual results of text-to-image is much worse than mask-to-image. To address this, we designed a uni-modal training mechanism with modal surrogates, yielding impressive results in multi-modal, uni-modal, and unconditional synthesis as shown in the third row from the top of Table  2. This approach leverages information from each modality and uses modal surrogates to learn inter-modal interactions, requiring only uni-modal data.
Entropy-aware modal-adaptive modulation. Enhancing our uni-modal training with modal surrogates, we integrated entropy-aware modal-adaptive modulation to adjust noise before it is altered by a fixed control strength. As illustrated in Table  2 and Figure  8, this method consistently outperforms others in multi-modal, uni-modal, and unconditional synthesis, especially in terms of fidelity to mask and text, while retaining high quality. Our entropy-aware approach ensures that all multi-modal conditions are given due consideration, leading to higher-fidelity image synthesis.
4.4 Extension to More Modalities
Our method’s notable flexibility and scalability enables effective learning of facial image generation from uni-modal training with modal surrogates. Additionally, our entropy-aware, modality-adaptive modulation fine-tunes denoising based on each modality’s specifics, ensuring harmonious multi-modality contributions. This versatility has led to the seamless integration of various controls such as sketches, low-resolution images, and 3D face models, enhancing our network’s capability for diverse facial synthesis tasks.
Diverse uni-modal synthesis. Figure  9 (a) demonstrates our model’s adeptness in uni-modal synthesis, consistently producing high-quality facial images with great fidelity to each modality. Whether it’s face shape from mask, an expression from a 3D face model, or details in a sketch, our single, versatile model captures these nuances effectively, showcasing its capability to maintain diversity while accurately reflecting each modality’s unique features.
Any combination of modal synthesis. Our method’s integration of various modalities allows for versatile combinations in face synthesis. For example, we can blend mask with lighting (Figure 9 (b)) or create faces based on poses and text (Figure 9 (c)). Our approach aids in text-guided super-resolution (Figure 9 (d)), and combines sketches with lighting effects (Figure 9 (e)). It can even merge lighting, text, sketch, and mask inputs to produce diverse, high-quality faces (Figure 9 (f)), demonstrating our method’s impressive capability in facial image synthesis.
@@ 5 Conclusion
In conclusion, our research presents a significant advancement in the field of multimodal face synthesis, presenting a highly scalable framework that supports face synthesis of high quality and fidelity under flexible combination of condition. Our approach introduces a uni-modal training framework with modal surrogates for each modality that serve as condition decorator for its modality and an inter-linker to facilitate inter-modal collaboration. The entropy-aware modal-adaptive modulation precisely tunes diffusion noise based on modal characteristics and conditions, enhancing the denoising process for superior synthesis quality. Our method significantly broadens multi-modal face synthesis capabilities, supporting a wide range of synthesis tasks from uni-modal to complex multi-modal combinations.

### 21 Mar 2024 v2

@@ 3.2 Entropy-Aware Modal-Adaptive Modulation
Multi-modal conditions naturally exhibits diverse entropy. Therefore it is essential to finely adjust the de-noise strategy according to multi-modal condition and modal-specific characteristic, ensuring generating vivid and faithful facial images. For uni-modal synthesis, one can flexibly specify the controlling strength to achieve high-quality synthesis results. However, in the multi-modal synthesis scenario, it is not desirable to employ a fixed controlling strength across all modalities. The reason behind this is that for some low controlling strength, the conditions of high condition entropy may have insufficient influences on image synthesis process. Simply increasing controlling strength may lead to overfitting of the generated images to some other condition of low condition entropy. We devise an entropy-aware modal-adaptive modulation to adaptively adjust the predicted noise according to given conditions and current de-noising state. As revealed in Fig. 5, we use a weighting module that processes the fused features $C_{m}$ of conditions and modal surrogates, and feature $f_{u}$ of the final decoding layer in U-Nnet. This network performs average pooling on these features, and then passes them through several linear layers to produce a weighting vector $w$. Note that the modal surrogates encapsulate modality-specific priors, and therefore they offer critical guidance on noise adjustment when combined with the given conditions. The inclusion of U-Net features also injects real-time insights $f_{u}$ into the de-noising trajectory. Apart from predicting a single base noise map based on the U-Net feature $f_{u}$, we follow the philosophy of multi-head [33] and predict multiple noise maps, each potentially capturing different aspects of the features. Finally, we can obtain the output noise that can be presented as:
$\begin{split}\epsilon_{\theta}=\frac{1}{K}\sum_{k}(w_{k}(n_{k}-n_{b})+n_{b}),\end{split}$ (6)
[…]
4 Experiments
In this section we first introduce our experimental settings in Sec. 4.1. Then we show various face synthesis applications based on our framework in Sec. 4.2. Next we conduct comparative study over existing competitive methods in Sec. 4.3. Finally we demonstrate effectiveness of our method via ablation analysis in Sec. 4.4.
4.1 Experimental Setup
Datasets. We conduct experiments on the Celeb-HQ dataset [15] with 30k high-quality face images. We utilized mask from [19] and textual descriptions from [14]. We compared our method, focusing on parsing mask and text modalities, for fair comparison with existing multi-modal methods following TediGAN [37] and collaborative Diffusion[13]. We train our method on the first 27K images and Our approach further efficiently includes additional modalities such as sketches [37], 3D face models [7], and low-resolution images [34], enriching our range of synthesis tasks.
Implementation details. We implement our method based on Latent Diffusion [29], where the images $(3\times H\times W)$are first mapped to latent features $(3\times\frac{H}{8}\times\frac{W}{8})$. The modal encoder of the low-resolution image is the same as the first stage encoder of LDM [29]. The encoder of text and low-resolution image is fixed during training, while the encoders of other modalities are updated in training. The modal surrogate for text is initialized as the textual encoder feature of input human face and others are randomly initialized. We train our method for 40k iterations on four RTX3090 GPUs with batch size of 32. We shall release our code upon publication of this manuscript.
4.2 Face Synthesis with Flexible Modal Combinations
Our uni-modal training strategy with modal surrogates enables us to involve more modalities without considering their complex combination, and linear increase of inference complexity. The entropy-aware modal-adaptive modulation further finely adjust noise according to given condition and modal-specific information, enabling various adaptive face synthesis. Therefore we could efficiently integrate more modalities and achieve a wide range of face synthesis applications, shown in Fig. 1 and Fig. 6.
Figure 7: Our method’s potential in integration with existing editing method[16] to achieve flexible editing tasks.
Diverse uni-modal synthesis. Fig.  6 (a) (left part) demonstrates our model’s adeptness in uni-modal synthesis, consistently producing high-quality facial images with great fidelity to various single modal condition. Whether the condition is face layout from mask, an expression from a 3D face model, or detailed structure in a sketch, our model captures these modal intrinsic accordingly and synthesis facial results of high fidelity, quality and diversity, within a single network. The pleasing uni-modal results demonstrate that our modal surrogate learns its modal intrinsic via Eq. 4 and decorate condition with modal-specific information, therefore our model can discriminate various types of conditions and synthesis face accordingly.
Diverse combination of multi-modal synthesis. Our method’s integration of various modalities allows for versatile combinations in face synthesis. For example, we can create face according mask layout with lighting effect (Fig. 6 (b)) or based on poses and text (Fig. 6 (c)). Our approach achieve text-guided super-resolution (Fig.  6 (d)), and combines sketches with lighting effects to generate face (Fig. 6 (e)). It can even merge lighting, text, sketch, and mask inputs to produce diverse, high-quality faces (Fig. 6 (f)), demonstrating our method’s impressive flexibility and adaptivity in various facial image synthesis tasks. Note that the synthesis results are obtained within a single sampling process of a single synthesis network regardless of the number of given conditions. We achieve face synthesis under flexible combination of multi-modal, no need to consider such complex modal combination during training. Instead we train our model in a uni-modal manner and our modal surrogate efficiently learns interaction among multi-modal conditions via Eq. 5.
Flexible multi-modal face editing. Fig. 7 demonstrates our method’s potential of integrating existing editing methods to achieve flexible editing tasks. We have incorporated editing method Imagic [16] into our multi-modal face synthesis framework. We begin by fixing the network parameters and optimizing the conditional embedding, which is initially set as the target condition feature. Next, the conditional is frozen and the diffusion U-Net parameters are further fine-tuned to reconstruct the image. The final step involves blending the feature embedding with the target embedding, guiding the diffusion network to generate the final results.
Method Training Data Type # of Synthesis Network FID $\downarrow$ Text(%) $\uparrow$ Mask (%) $\uparrow$
TediGAN [37] Uni-modal 1 116.04 24.38 86.37
Compose [21, 23] Uni-modal 2 127.39 24.22 77.34
CoDiff [13] Multi-modal 2 122.51 24.37 85.94
ControlNet [40] Multi-modal 1 136.41 25.70 85.44
T2I-Adapter [22] Multi-modal 1 139.82 26.11 79.84
Ours Uni-modal 1 103.14 24.70 90.16
Mask
Text
TediGAN
Compose
CoDiff
T2IAdapter
ControlNet
The woman looks young. No beard on her face
He is a teenager. His face is covered with his stubble.
Mask
Text
TediGAN
Compose
CoDiff
T2IAdapter
ControlNet
@@ 4.3 Comparison Analysis
Quantitative and qualitative comparisons. In Tab. 1, we benchmark our method’s FID, Text, and Mask accuracy in multi-modal facial synthesis against leading techniques. TediGAN[37] yields high quality results by optimizing each condition and using a high-resolution generator StyleGAN. As shown in Fig. 8, however, TediGAN struggles in producing results of satisfying alignment and it involves per-instance fine-tuning for each condition. The generated face either overlooks the mask detail (hair shape of woman) or neglects teenager from text. ControlNet [40] and T2I-Adpter [22] score well in text accuracy, leveraging advanced text encoders[27] and inheriting capabilities from Stable Diffusion[29]. As shown in Fig. 8 and Tab. 1, this line of methods requires multi-modal annotated data and produce noticeable artifacts. Compose[23, 21] lags behind due to its simplistic fusion of uni-modal model noise, overlooking crucial modality-specific and inter-modal collaboration in multi-modal synthesis. The generated face exhibits notable artifacts. Our method produce results of high quality that is align well with given conditions, e.g. the hair detail of given mask and the age and beard attribute from given text. We achieve multi-modal face synthesis within a single synthesis network and require uni-modal annotate data under efficient uni-modal training.
User study. We have invited 12 experienced researchers to score the results of compared methods, in terms of image quality and condition alignment. The higher the score, the better the results. Fig. 9 shows that our method yields better performance in terms of image quality and condition alignment.
Figure 9: User study. Our results achieve best among compared methods in terms with fidelity and quality of generated face.
@@ 4.4 Ablation Analysis
We first validate how our uni-modal training with surrogate helps to achieve scalable and flexible multi-modal face synthesis. Next, we examine the effectiveness of our entropy-aware modal-adaptive modulation in fully leveraging the control of varing multi-modal conditions.
Modal surrogate function: condition decoration. We demonstrate the effectiveness of modal surrogate to decorate condition. As shown in Tab. 2, the baseline $M_{1}$ first trains the parallel synthesis network for each modality and then fuses noise of each modal to achieve multi-modal face synthesis. $M_{1}$ produces acceptable results for the two uni-modal task (text/mask to face), but suffers from poor scalability for its linear increase for the parameter number and sampling complexity of the modal number. $M_{2}$ improves $M_{1}$ by setting for each modality a trainable modal surrogate to decorate condition of its modal. The resulting network is capable of generating face under various type of conditions within a single model.
Table 2: Ablation results of uni-modal training with modal surrogates and entropy-aware modal-adaptive modulation. For each task the top three are marked in red, blue, and green, respectively.
$M_{1}$ $M_{2}$ $M_{3}$ $M_{4}$ $M_{5}$ $M_{6}$
# of Model Parallel model Shared model Shared model Shared model Shared model Shared model
Training strategy Uni-modal train Uni-modal train Uni-modal train Multi-modal train Multi-modal train Uni-modal train
Surrogate - Condition decoration Condition decoration - Condition decoration Condition decoration
function - - Inter-modal learning - - Inter-modal learning
Adjust Noise - - - - -
[…]
Mask to Face 113.77 - 87.21 109.31 - 89.27 109.83 - 88.97 - - - 108.69 - 90.06 112.91 - 90.31
Text to Face 108.81 24.54 - 109.61 24.29 - 114.07 24.48 - - - - 195.94 23.18 - 110.32 24.78 -
Mask&Text to Face 117.94 24.24 78.55 121.77 24.07 88.33 113.29 24.28 89.72 124.98 24.28 85.44 112.25 24.42 89.91 103.14 24.70 90.16
Figure 10: Comparison of synthesis results across training methodologies. Our uni-modal training with modal surrogate enables a flexible and scalable face synthesis framework. Entropy-aware Modality-Adaptive Modulation further enhances the fidelity to given conditions.
Figure 11: Further verification of our method’s effectiveness to learn inter-modal collaboration for multi-modal synthesis. T2I-Adapter and Uni-control directly combine features from multi-modal layout, and they do not modelling inter-modal collaboration.
Modal surrogate function: inter-modal learning. We demonstrate the effectiveness of modal surrogate to learn inter-modal interaction in helping achieve multi-modal face synthesis. Both $M_{1}$ and $M_{2}$ produce poor results for multi-modal synthesis, shown in Tab. 2 and Fig. 10. That is because their training process do not involve how multi-modal conditions interact in face synthesis process. $M_{3}$ improves $M_{2}$ by extra involving all other modal surrogates during training. In this way the surrogates of some modal learned from data of other modalities, and thus serve as inter-modal linker for multi-modal synthesis. Tab. 2 and Fig. 10 demonstrate that compared with $M_{2}$, $M_{3}$ supports multi-modal synthesis (text & mask to face) of higher quality and condition alignment. We further compare the method of combining uni-modal noise or features and our method in multi-modal synthesis in Fig. 11. Directly combining uni-modal feature or noise tends to yield poor results as it lacks of multi-modal modelling in multi-modal synthesis. In contrast, our method effectively fuses constraints from multiple layouts. Our cross-modal updating of surrogates helps to grasp multi-modal interaction with efficient uni-modal training.
Uni-modal Training Our uni-modal training helps network avoid the short cut issue in multi-modal learning, and fully learns synthesis for each modality. $M_{5}$ represents the method of multi-modal training with modal surrogate as condition decoration. Tab. 2 and Fig. 10 demonstrate that $M_{5}$ tends to rely more on mask and overlook text condition, as the performance of text synthesis is rather bad. In contrast our uni-modal training strategy $M_{3}$ fully learns synthesis under each modality condition, thus resulting flexible synthesis versatile across all modalities.
Entropy-aware modal-adaptive modulation. Building upon uni-modal training with modal surrogate $M_{3}$, we further integrate entropy-aware modal-adaptive modulation $M_{6}$ to adjust noise level in accordance with the given condition. Fig. 10 shows that $M_{3}$ neglect some constraint of multi-modal condition, e.g., the man face has no beard. $M_{3}$ tends to overlook conditions of high entropy (text) for multi-modal synthesis. $M_{6}$ shows improvement in terms with condition alignment of both conditions for both uni-modal and multi-modal synthesis. In the right part of Fig. 12 (w/o EAM) shows that even finely adjusting the user-defined guidance weight $w$ of CFG [11] still fails to obtain results of satisfying quality and alignment. In contrast, our full method (with EAM) exhibits high fidelity and image quality due to our finely, real-time adjustment of the de-noising level based on the conditions and the current state of de-noising.
@@ 5 Conclusion
In conclusion, our research presents a significant advancement in the field of multi-modal face synthesis, presenting a highly scalable framework that supports face synthesis of high quality and fidelity under flexible combination of condition. Our approach introduces a uni-modal training framework with modal surrogates for each modality that serve as condition decorator for its modality and an inter-linker to facilitate inter-modal collaboration. The entropy-aware modal-adaptive modulation precisely tunes diffusion noise based on modal characteristics and conditions, enhancing the de-noising process for superior synthesis quality. Our method broadens multi-modal face synthesis capabilities, supporting a wide range of synthesis tasks from uni-modal to complex multi-modal combinations.